# MedQARo: A Large-Scale Benchmark for Medical Question Answering in Romanian

### v2

@@ 3.2 Preprocessing
To support model training and evaluation, we constructed two domain-specific datasets targeting question answering in oncology given a clinical context: one for breast cancer and one for lung cancer.
@@
3.3 Data Partitioning and Statistics
We split the dataset into training, validation and test, using a $70\%/15\%/15\%$ split ratio, ensuring that the division occurs at the patient level. We prefer this split over splitting individual QA pairs because it guarantees that all the data from a single patient, including their epicrisis and associated QA pairs, remains within exactly one subset, thereby preventing data leakage among splits.
We present the distribution of our dataset across training, validation, and test sets in Table 2. For the breast cancer subset, our split results in 557 patients allocated for training, 119 for validation, and 120 for testing. In terms of QA pairs, the split distributes 53,472 samples for training, 11,424 for validation, and 11,520 for testing, respectively. Similarly, the lung cancer subset is divided into 150 training, 32 validation, and 33 test patients, yielding 18,300, 3,904, and 4,026 QA pairs, respectively. The splits are independently generated for each cancer type. We next combine the corresponding partitions to create our final unified dataset, called MedQARo. In total, MedQARo comprises 1,011 unique patients (796 suffering from breast cancer, and 215 from lung cancer) and 102,646 question-answer pairs, providing a substantial resource for training and evaluating Romanian medical QA systems.
In Table 3, we report statistics about the number of tokens in epicrises, questions and answers included MedQARo. Epicrises range from 207 to 34,247 tokens, with a mean value of 7,171 tokens, where shorter documents typically correspond to patients under investigation or at early treatment stages, while longer ones capture extensive treatment histories with multiple interventions. This variation can pose challenges for some models with limited context windows. While epicrises are typically lengthy, questions and answers are much shorter. Questions maintain relative consistency, between 7 and 30 tokens, with a mean value of 15 tokens, reflecting the consistency of clinical query formulations. The answers are represented by predominantly brief texts with a mean of 4 tokens, varying in length from one to 442 tokens. Longer responses are typical for questions involving complex medical reasoning.
@@ 3.3 Data Partitioning and Statistics
Epicrisis 207 34,247 7,171 7,249,881
Question 7 30 15 1,621,561
Answer 1 442 4 420,708
@@
We categorize the questions into three types: (i) binary (yes/no); (ii) extractive, where the answer appears explicitly in the clinical context (epicrisis); and (iii) reasoning, where the answer must be inferred from multiple clues within the epicrisis. The distribution of these categories in MedQARo is shown in Figure 2. Within the binary subset, the label distribution is moderately skewed, with roughly two thirds of answers being “Nu” (“No”) and one third being “Da” (“Yes”). The slightly skewed distribution resulted naturally from the data collection process.
@@ 4 Methods
To address the task of medical question answering in Romanian, we consider a diverse set of LLMs and use them under various configurations, considering both zero-shot and fine-tuning settings. Our goal is to understand how distinct architectural specializations (e.g. language-specific, domain-specific, or long-context models) and alternative prompt formulations affect performance on the Romanian clinical QA task. We select four distinct LLM families. Two of them are pretrained on Romanian, namely RoLLaMA2-7B Masala et al. (2024), which is adapted from LLaMA 2 Touvron et al. (2023), and RoMistral-7B Masala et al. (2024), which is adapted from the Mistral architecture Jiang et al. (2023). The third model, namely Phi-4-mini-instruct Abouelenin et al. (2025), is optimized for long-context inputs. This is useful for the processing of entire epicrises, which can often get to be very long (see Table 3), especially for patients with a long medical history. Finally, given the target domain, the fourth model, LLaMA3-OpenBioLLM-8B Pal and Sankarasubbu (2024), is pretrained on biomedical corpora.
All four LLMs are employed in two setups, zero-shot prompting and end-to-end fine-tuning. Due to the large number of parameters in the chosen models, we adopt a parameter-efficient fine-tuning strategy based on Low-Rank Adaptation (LoRA) Hu et al. (2022) specialized on the causal language modeling task. By comparing the results obtained in the two setups, we can precisely contextualize the impact of fine-tuning on MedQARo. In addition, we also test two alternative prompts, where the order between context (epicrisis) and question is interchanged. Finally, to determine the impact of zero-shot prompting, we include a random token selector and a majority answer model as baselines for the QA task. Next, we describe all evaluated methods in detail.
Figure 2: Distribution of question types in the MedQARo dataset. The dataset comprises three main categories: binary (yes/no) questions, extractive questions (answers explicitly found in the epicrisis), and reasoning questions (requiring inference beyond explicit mentions). Percentages are computed over the full set of 102,646 QA pairs. Best viewed in color.
@@ 4.1 Romanian-Adapted LLMs
We start from two instruction-tuned language models specifically adapted for Romanian, RoLLaMA2-7B-Instruct and RoMistral-7B-Instruct. Both models are part of the recent efforts to build high-quality Romanian LLMs by further pretraining on Romanian corpora and applying instruction tuning with Romanian datasets, such as RoAlpaca Taori et al. (2023); Masala et al. (2024), RoOpenAssistant Köpf et al. (2023); Masala et al. (2024) and RoOrca Mukherjee et al. (2023); Masala et al. (2024). The two LLMs have been optimized for Romanian morphology and syntax using tokenizers adapted to the language. They have shown strong performance on Romanian natural language understanding and generation tasks Masala et al. (2024). In our study, we further fine-tune both models on MedQARo and compare their performance under different prompt structures and lengths.
RoLLaMA2-7B-Instruct. As it inherits the architectural constraints of the original LLaMA 2 model Touvron et al. (2023), RoLLaMA2-7B is limited to a context window of 4,096 tokens. We employ a LoRA configuration specialized on the causal language modeling task. The chosen configuration enables updating a small subset of the model parameters ($0.0622\%$) during fine-tuning, significantly reducing training overhead. To enhance the adaptability of the LLM to the clinical QA task, we further consider several configuration alternatives. First, we evaluate on various prompt lengths, considering limits of 1,024, 2,048 and 4,096 tokens. If an epicrisis is longer than the accepted input length, we simply trim it from the end. When evaluating on prompts of 4,096 tokens, we also explore two prompt formats: Q+E+A (question + epicrisis + answer) and E+Q+A (epicrisis + question + answer). The model is trained via token-level cross-entropy, where the task is to predict the answer tokens via next token prediction. For this reason, the answer is always placed at the end of the prompt.
RoMistral-7B-Instruct. Given the large number of parameters in the base Mistral model, we again employ a LoRA-based fine-tuning strategy, enabling updates to only a small number of parameters. In this case, we fine-tune just $0.0470\%$ of the parameters, using the same LoRA configuration as for RoLLaMA2-7B-Instruct. This approach ensures both memory efficiency and effective domain adaptation to the Romanian clinical QA setting. We conduct experiments with two prompt lengths, namely 2,048 and 4,096. Consistent with the preliminary empirical results obtained with the previous model, the prompt format is set to Q+E+A in all the experiments performed with RoMistral-7B-Instruct. We employ the token-level cross-entropy to optimize the model.
@@ 4.2 Long-Context LLM
To adapt the model to the clinical QA setting, we fine-tune it using LoRA, updating only $0.0956\%$ of the parameters via token-level cross-entropy. We employ the same LoRA configuration as for RoLLaMA2-7B-Instruct and RoMistral-7B-Instruct. Thanks to the architectural support of Phi-4 for extended input lengths, we evaluate the model under increasingly longer prompts, namely 2,048, 3,072, 4,096, 8,192, and 16,384 tokens. This allows us to assess the relationship between available context and model performance, particularly in scenarios where clinical answers are distributed across epicrises of patients with a long medical history. All prompts follow the Q+E+A format.
@@ 4.4 Baselines
To assess the effectiveness of our fine-tuned models and to better isolate the impact of domain and task adaptation, we include three baselines.
@@
Model Prompt format Validation Test F1 EM BLEU METEOR F1 EM BLEU METEOR RoLLaMA2-7B-Instruct E+Q+A 0.2518 0.2395 0.2233 0.1656 0.2871 0.2800 0.2531 0.1867 Q+E+A 0.2763 0.2545 0.2508 0.1452 0.3142 0.2890 0.2772 0.1647
@@ 5 Experiments
We conduct experiments with the chosen methods on the MedQARo dataset, using our official split. We start by tuning various hyperparameters involved in the fine-tuning process, aiming to find the optimal hyperparameter setup for all LLMs. We next evaluate two different prompt formats, to determine which prompt format leads to higher performance. For the best performing prompt structure, we conduct several experiments across multiple LLMs, using various prompt lengths. The objectives of our experiments are: (i) to capture how Romanian-adapted models compare to general-purpose and domain-specific alternatives, (ii) to determine how fine-tuning performs against zero-shot inference, (iii) to establish how fine-tuned or zero-shot LLMs compare against trivial baselines, and (iv) to investigate how the context length influences model performance, particularly for models with extended context capabilities.
@@ 5.1 Evaluation
BLEU. Because MedQARo contains answers with more than one token, we consider the BLEU score Papineni et al. (2002), which measures n-gram overlap to assess fluency and lexical similarity. When the predicted answer includes only a subset of the tokens from the ground-truth answer, BLEU can provide a representative measure by considering partial n-gram matches.
METEOR. Lastly, we include METEOR Banerjee and Lavie (2005), which offers a more flexible evaluation by accounting for synonyms, stemming, and word order variations. These aspects are particularly valuable given the rich morphology and flexible syntax of Romanian.
Model Fine-tuned #Tokens Validation Test F1 EM BLEU METEOR F1 EM BLEU METEOR Random token selector N/A 1,024 0.0023 0.0006 0.0006 0.0013 0.0019 0.0001 0.0004 0.0011 2,048 0.0021 0.0004 0.0004 0.0013 0.0024 0.0003 0.0005 0.0015 4,096 0.0022 0.0005 0.0006 0.0014 0.0026 0.0003 0.0005 0.0015 Majority answer N/A N/A 0.2091 0.1786 0.1831 0.1153 0.2148 0.1838 0.1884 0.1183 RoLLaMA2-7B-Instruct 2,048 0.0188 0.0015 0.0024 0.0146 0.0230 0.0025 0.0023 0.0166 1,024 0.5535 0.5431 0.4996 0.3498 0.5646 0.5300 0.4896 0.3451 2,048 0.5832 0.6035 0.5651 0.3487 0.5944 0.6145 0.5780 0.3554 4,096 0.2763 0.2545 0.2508 0.1452 0.3142 0.2890 0.2772 0.1647 RoMistral-7B-Instruct 4,096 0.0715 0.0430 0.0234 0.0438 0.0605 0.0350 0.0229 0.0403 2,048 0.5761 0.5820 0.5387 0.3425 0.5808 0.5800 0.5331 0.3477 4,096 0.5814 0.5925 0.5487 0.3456 0.5963 0.6045 0.5546 0.3590 Phi-4-mini-instruct 2,048 0.0078 0.0000 0.0007 0.0075 0.0048 0.0002 0.0005 0.0043 2,048 0.6500 0.6846 0.6366 0.3916 0.6618 0.6935 0.6468 0.3976 3,072 0.6617 0.6960 0.6472 0.3974 0.6671 0.7011 0.6522 0.4009 4,096 0.4217 0.4145 0.3923 0.2274 0.4220 0.4155 0.3799 0.2303 8,192 0.5050 0.4960 0.4624 0.2911 0.5216 0.5105 0.4653 0.3037 16,384 0.5168 0.5110 0.4765 0.2991 0.5267 0.5130 0.4731 0.3060 LLaMA3-OpenBioLLM-8B 2,048 0.0140 0.0004 0.0014 0.0156 0.0119 0.0004 0.0015 0.0132 2,048 0.5861 0.6140 0.5720 0.3498 0.5958 0.6215 0.5792 0.3585 4,096 0.5365 0.5855 0.5210 0.3163 0.5566 0.5775 0.5382 0.3348
Table 5: Results of multiple LLMs based on various configurations vs. two trivial baselines on MedQARo. For each LLM, there is a zero-shot version and several fine-tuned versions, with various prompt lengths. F1, Exact Match (EM), BLEU, and METEOR scores are reported on both validation and test sets. Fine-tuned models consistently outperform zero-shot versions. The best results for each LLM are highlighted in bold green.

### v3

@@ 3.2 Preprocessing
To support model training and evaluation, we constructed two domain-specific datasets targeting question answering in oncology given a clinical context: one for breast cancer and one for lung cancer. These are complemented by a generic dataset targeting generic questions covering five other cancer types (not including breast and lung cancer).
@@
Other cancers. To construct a benchmark that enables the assessment of generalization capacity across medical centers and cancer types, we created a secondary dataset, which is not used during training. This dataset comprises 231 patients, diagnosed with a different kind of cancer (head and neck, melanoma, renal cell carcinoma, bladder cancer, or hepatocellular carcinoma), all coming from a different center than the patients included in our main dataset. For these patients, there are 7 general questions, relevant for all five cancer types. As for breast and lung cancer patients, we randomly sampled two different question variants for each patient-question pair, leading to a total of 3,234 unique QA pairs.
3.3 Partitioning and Statistics
We merge the breast and lung cancer patients (from the Colţea Clinical Hospital) and corresponding QA pairs into a single dataset. Then, the merged dataset is split into training, validation and in-domain test, using a $70\%/15\%/15\%$ split ratio, ensuring that the division occurs at the patient level. We prefer this split over splitting individual QA pairs because it guarantees that all the data from a single patient, including their epicrisis and associated QA pairs, remains within exactly one subset, thereby preventing data leakage among splits. Furthermore, we keep the patients with other cancers (from the “Dan Furtună” Medical Center) for cross-domain evaluation.
@@ 3.3 Partitioning and Statistics
Epicrisis 207 34,247 7,829 9,723,618
Question 7 30 15 1,664,296
Answer 1 442 4 432,754
Figure 2: Distribution of question types in the MedQARo dataset. The dataset comprises three main categories: binary (yes/no) questions, extractive questions (answers explicitly found in the epicrisis), and reasoning questions (requiring inference beyond explicit mentions). Percentages are computed over the full set of 105,880 QA pairs. Best viewed in color.
We present the distribution of our dataset across training, validation, and test sets in Table 2. For the breast cancer subset, our split results in 557 patients allocated for training, 119 for validation, and 120 for testing. In terms of QA pairs, the split distributes 53,472 samples for training, 11,424 for validation, and 11,520 for testing, respectively. Similarly, the lung cancer subset is divided into 150 training, 32 validation, and 33 test patients, yielding 18,300, 3,904, and 4,026 QA pairs, respectively. A separate cross-domain test set contains 231 patients and 3,234 QA pairs. In summary, MedQARo contains four data partitions: one for training (707 patients), one for in-domain validation (151 patients), one for in-domain testing (153 patients), and one for cross-domain testing (231 patients). In total, MedQARo comprises 1,242 unique patients (796 suffering from breast cancer, 215 from lung cancer, and 231 from five other cancers) and 105,880 question-answer pairs. Hence, our contribution is a substantial resource for training and evaluating Romanian medical QA systems, in both in-domain and cross-domain scenarios.
In Table 3, we report statistics about the number of tokens in epicrises, questions and answers included in MedQARo. Epicrises range from 207 to 34,247 tokens, with a mean value of 7,829 tokens, where shorter documents typically correspond to patients under investigation or at early treatment stages, while longer ones capture extensive treatment histories with multiple interventions. This variation can pose challenges for some models with limited context windows. While epicrises are typically lengthy, questions and answers are much shorter. Questions maintain relative consistency, between 7 and 30 tokens, with a mean value of 15 tokens, reflecting the consistency of clinical query formulations. The answers are represented by predominantly brief texts with a mean of 4 tokens, varying in length from one to 442 tokens. Longer responses are typical for questions involving complex medical reasoning.
Table 4: Representative examples of binary, extractive, and reasoning questions from MedQARo. Each question is shown in Romanian, followed by its English translation. One reformulation is displayed below each distinct question.
Type Question (Romanian / English)
Binary S-a realizat o intervenţie chirurgicală înainte de iniţierea imunoterapiei? / Has a surgical intervention been performed before initiating immunotherapy?
Intervenţia chirurgicală a precedat imunoterapia în cazul acestui pacient? / Did surgery precede immunotherapy for this patient?
Pacientul are gena BRCA1? / Does the patient have the BRCA1 gene?
Rezultatul analizei genetice indică prezenţa mutaţiei BRCA1? / Does the genetic analysis result indicate the presence of the BRCA1 mutation?
S-au identificat metastaze la nivelul plămânilor? / Have lung metastases been identified?
Există determinări secundare pulmonare la acest pacient? / Are there any secondary pulmonary tumors in this patient?
Extractive Care a fost dimensiunea tumorii la debut? / What was the tumor size at the time of diagnosis?
Cât de mare era tumora atunci când a fost stabilit diagnosticul? / How large was the tumor when the diagnosis was made?
Ce cadrane ale sânului sunt invadate de tumoră? / Which breast quadrants are invaded by the tumor?
Specifică cadranele invadate de tumoră / Specify the quadrants invaded by the tumor
Ce status ECOG are pacientul? / What is the ECOG status of the pacient?
Specifică statusul de performanţă ECOG / Specify the ECOG performance status
Reasoning Câte luni sunt de la începutul tratamentului până la apariţia progresiei bolii? / How many months have passed from the start of treatment until the onset of disease progression?
Cât a durat (în luni) ca boala să progreseze sub tratament? / How long did it take (in months) for the disease to progress under treatment?
Câte luni au trecut de la chirurgie până la începerea tratamentului cu radioterapie? / How many months have elapsed from surgery until the initiation of radiotherapy treatment?
La cât timp după operaţie a început tratamentul radioterapic? / How long after surgery did radiotherapy treatment begin?
Care este stadiul tumoral la diagnostic al pacientului? / What was the tumor stage at the time of diagnosis for this patient?
Ce stadiu de cancer a fost stabilit la diagnostic? / What stage of cancer was determined at diagnosis?
(a) Age distribution for breast cancer.
(b) Gender distribution for breast cancer.
(c) Age distribution for lung cancer.
(d) Gender distribution for lung cancer.
(e) Age distribution for other cancers.
(f) Gender distribution for other cancers.
Figure 3: Age and gender distributions of patients included in MedQARo. Rows correspond to breast cancer, lung cancer, and other cancers, respectively. The first column shows age distributions using 5-year bins (between 20 and 94 years), while the second column row shows gender distributions. Best viewed in color.
@@
We categorize the questions into three types: (i) binary (yes/no); (ii) extractive, where the answer appears explicitly in the clinical context (epicrisis); and (iii) reasoning, where the answer must be inferred from multiple clues within the epicrisis. The distribution of these categories in MedQARo is shown in Figure 2. Within the binary subset, the label distribution is moderately skewed, with roughly two thirds of answers being “Nu” (“No”) and one third being “Da” (“Yes”). The slightly skewed distribution resulted naturally from the data collection process. In Table 4, we showcase some representative examples from the three question categories included in MedQARo. Some questions are specific to a certain cancer type, e.g. “Ce cadrane ale sânului sunt invadate de tumoră?” / “Which breast quadrants are invaded by the tumor?” is only relevant for breast cancer, while other questions are valid for all cancer types, e.g. “Care este stadiul tumoral la diagnostic al pacientului?” / “What was the tumor stage at the time of diagnosis for this patient?” is applicable to all cancers. The reformulated questions indicate that semantic equivalence was ensured during the process of manual question generation conducted by the annotators.
3.4 Demographics
We further examine the age and gender distributions of patients included in MedQARo. In Figure 3, we report the distributions for the three patient groups, which are separated by cancer type: breast cancer, lung cancer, and other cancers. Across all diagnostic categories, the cohort is predominantly composed of middle-aged and elderly patients, with most cases falling in the 50-79 age range. As naturally expected, most breast cancer patients are females (793 out of 796), as breast cancer is less common in the male population. In contrast, the male population is more affected by lung cancer and other cancers, which can be attributed to the often negligent behavior of males towards maintaining a healthy lifestyle [Bonhomme-JMHG-2007, Dong-G-2012].
In Figure 4, we illustrate the distribution of patients diagnosed with other cancer types. The distribution reveals heterogeneous demographic patterns across diagnoses, with head and neck cancers accounting for the largest proportion of cases, followed by melanoma.
Figure 4: Distribution across cancer types for patients included in the cross-domain test set of MedQARo. Best viewed in color.
These demographic patterns are consistent with real-world clinical populations and suggest that MedQARo captures representative patient characteristics, supporting the etiological validity of the benchmark for clinical question answering.
@@ 4 Methods
To address the task of medical question answering in Romanian, we consider a diverse set of LLMs and use them under various configurations, considering both zero-shot and fine-tuning settings. Our goal is to understand how distinct architectural specializations (e.g. language-specific, domain-specific, or long-context models) and alternative prompt formulations affect performance on the Romanian clinical QA task. We select six distinct LLM families. Two of them are pretrained on Romanian, namely RoLLaMA2-7B [masala2024vorbecstiromanecsterecipetrain], which is adapted from LLaMA 2 [touvron2023llama], and RoMistral-7B [masala2024vorbecstiromanecsterecipetrain], which is adapted from the Mistral architecture [jiang2023mistral7b]. The third model, namely Phi-4-mini-instruct [phi4mini], is optimized for long-context inputs. This is useful for the processing of entire epicrises, which can often get to be very long (see Table 3), especially for patients with a long medical history. Given the target domain, the fourth model, LLaMA3-OpenBioLLM-8B [OpenBioLLMs], is pretrained on biomedical corpora. Finally, we evaluate two state-of-the-art LLMs, GPT-5.2 [GPT52-2025] and Gemini 3 Flash [Gemini3-2025], that are only accessible via public APIs.
The first four LLMs are employed in two setups, zero-shot prompting and end-to-end fine-tuning. Due to the large number of parameters in the chosen models, we adopt a parameter-efficient fine-tuning strategy based on Low-Rank Adaptation (LoRA) [hu2021lora] specialized on the causal language modeling task. By comparing the results obtained in the two setups, we can precisely contextualize the impact of fine-tuning on MedQARo. In addition, we also test two alternative prompts, where the order between context (epicrisis) and question is interchanged. To determine the impact of zero-shot prompting, we include a random token selector and a majority answer model as baselines for the QA task. Since access to the architectures and weights of GPT-5.2 and Gemini 3 Flash is not available, we only employ these two models in the zero-shot setup. Next, we describe all evaluated methods in detail.
@@ 4.1 Romanian-Adapted LLMs
We start from two instruction-tuned language models specifically adapted for Romanian, RoLLaMA2-7B-Instruct and RoMistral-7B-Instruct. Both models are part of the recent efforts to build high-quality Romanian LLMs by further pretraining on Romanian corpora and applying instruction tuning with Romanian datasets, such as RoAlpaca [alpaca, masala2024vorbecstiromanecsterecipetrain], RoOpenAssistant [kopf2024openassistant, masala2024vorbecstiromanecsterecipetrain] and RoOrca [mukherjee2023orca, masala2024vorbecstiromanecsterecipetrain]. The two LLMs have been optimized for Romanian morphology and syntax using tokenizers adapted to the language. They have shown strong performance on Romanian natural language understanding and generation tasks [masala2024vorbecstiromanecsterecipetrain]. In our study, we further fine-tune both models on MedQARo and compare their performance under different prompt structures and lengths.
RoLLaMA2-7B-Instruct. As it inherits the architectural constraints of the original LLaMA 2 model [touvron2023llama], RoLLaMA2-7B is limited to a context window of 4,096 tokens. We employ a LoRA configuration specialized on the causal language modeling task. The chosen configuration enables updating a small subset of the model parameters ($0.0622\%$) during fine-tuning, significantly reducing training overhead. To enhance the adaptability of the LLM to the clinical QA task, we further consider several configuration alternatives. First, we evaluate on various prompt lengths, considering limits of 1,024, 2,048 and 4,096 tokens. If an epicrisis is longer than the accepted input length, we simply trim it from the end. When evaluating on prompts of 4,096 tokens, we also explore two prompt formats: Q+E+A (question + epicrisis + answer) and E+Q+A (epicrisis + question + answer). The model is trained via token-level cross-entropy, where the task is to predict the answer tokens via next token prediction. For this reason, the answer is always placed at the end of the prompt.
RoMistral-7B-Instruct. Given the large number of parameters in the base Mistral model, we again employ a LoRA-based fine-tuning strategy, enabling updates to only a small number of parameters. In this case, we fine-tune just $0.0470\%$ of the parameters, using the same LoRA configuration as for RoLLaMA2-7B-Instruct. This approach ensures both memory efficiency and effective domain adaptation to the Romanian clinical QA setting. We conduct experiments with two prompt lengths, namely 2,048 and 4,096. Consistent with the preliminary empirical results obtained with the previous model, the prompt format is set to Q+E+A in all the experiments performed with RoMistral-7B-Instruct. We employ token-level cross-entropy to optimize the model.
@@ 4.2 Long-Context LLM
To adapt the model to the clinical QA setting, we fine-tune it using LoRA, updating only $0.0956\%$ of the parameters via token-level cross-entropy. We employ the same LoRA configuration as for RoLLaMA2-7B-Instruct and RoMistral-7B-Instruct. Thanks to the architectural support of Phi-4 for extended input lengths, we evaluate the model under increasingly longer prompts, namely 2,048, 3,072, 4,096, 8,192, 16,384 and 32,768 tokens. This allows us to assess the relationship between available context and model performance, particularly in scenarios where clinical answers are distributed across epicrises of patients with a long medical history. All prompts follow the Q+E+A format.
@@ 4.4 Baselines
To assess the effectiveness of our fine-tuned models and to better isolate the impact of domain and task adaptation, we include four baselines.
@@
Zero-shot API-based LLMs. In addition to the open-source models, we evaluate two state-of-the-art LLMs that are only accessible through commercial APIs, namely GPT-5.2 and Gemini 3 Flash. GPT-5.2 is designed to effectively process long contexts, have strong reasoning capabilities, and support multiple languages. Its reasoning ability is enforced via reinforcement learning, while the multilingual support is ensured by the large-scale pretraining on public data. Gemini 3 Flash is an LLM constructed on top of Gemini 3 Pro, inheriting its multimodal and reasoning capabilities. Its architecture is based on a sparse mixture-of-experts (MoE), which is able to activate only a subset of parameters (experts) per input token. This is achieved by learning to dynamically route tokens to a subset of experts. The thinking levels of Gemini 3 Flash are configured to strike a balance between quality, cost and latency. For both models, we use the same prompt format as before, namely Q+E+A, to ensure a fair comparison with the other models. To control evaluation costs, we compare GPT-5.2 and Gemini 3 Flash with the other LLMs on a randomly sampled subset of 100 test instances. The comparative empirical study is intended to provide a reference point for the zero-shot capabilities of current state-of-the-art LLMs on Romanian clinical question answering.
@@ 5 Experiments
We conduct experiments with the chosen methods on the MedQARo dataset, using our official data split. We start by tuning various hyperparameters involved in the fine-tuning process, aiming to find the optimal hyperparameter setup for all LLMs. We next evaluate two different prompt formats, to determine which prompt format leads to higher performance. For the best performing prompt structure, we conduct several experiments across multiple LLMs, using various prompt lengths. The objectives of our experiments are: (i) to capture how Romanian-adapted models compare with general-purpose and domain-specific alternatives, (ii) to determine how fine-tuning performs against zero-shot inference, (iii) to establish how fine-tuned or zero-shot LLMs compare against trivial baselines, (iv) to investigate how the context length influences model performance, particularly for models with extended context capabilities, and (v) to determine the generalization capacity of LLMs across distinct data distributions, where patients come from different medical centers, and suffer from different cancer types.
@@ 5.1 Evaluation
BLEU. Because MedQARo contains answers with more than one token, we consider the BLEU score [papineni2002bleu], which measures n-gram overlap to assess fluency and lexical similarity. When the predicted answer includes only a subset of the tokens from the ground-truth answer, BLEU can provide a representative measure by considering partial n-gram matches.
METEOR. Lastly, we include METEOR [banerjee2005meteor], which offers a more flexible evaluation by accounting for synonyms, stemming, and word order variations. These aspects are particularly valuable given the rich morphology and flexible syntax of Romanian.
Model Prompt format Validation In-domain test F1 EM BLEU METEOR F1 EM BLEU METEOR RoLLaMA2-7B E+Q+A 0.2518 0.2395 0.2233 0.1656 0.2871 0.2800 0.2531 0.1867 Q+E+A 0.2763 0.2545 0.2508 0.1452 0.3142 0.2890 0.2772 0.1647